OpenStudy (dtan5457):

if f(ax)=af(x), f(15)=3, compute f(49)

OpenStudy (dtan5457):

@freckles

jimthompson5910 (jim_thompson5910):

This might work f(15)=3 f(15*1)=3 15*f(1)=3 f(1)=3/15 f(1)=1/5 49*f(1)=49*(1/5) f(49*1)=49/5 f(49)=49/5

OpenStudy (dtan5457):

easy and simple approach, thanks!